# ELiC: Efficient LiDAR Geometry Compression via Cross-Bit-depth Feature Propagation and Bag-of-Encoders

### v1

@@ Abstract
Hierarchical LiDAR geometry compression encodes voxel occupancies from low to high bit-depths, yet prior methods treat each depth independently and re-estimate local context from coordinates at every level, limiting compression efficiency. We present ELiC, a real-time framework that combines cross-bit-depth feature propagation, a Bag-of-Encoders (BoE) selection scheme, and a Morton-order-preserving hierarchy. Cross-bit-depth propagation reuses features extracted at denser, lower depths to support prediction at sparser, higher depths. BoE selects, per depth, the most suitable coding network from a small pool, adapting capacity to observed occupancy statistics without training a separate model for each level. The Morton hierarchy maintains global Z-order across depth transitions, eliminating per-level sorting and reducing latency. Together these components improve entropy modeling and computation efficiency, yielding state-of-the-art compression at real-time throughput on Ford and SemanticKITTI. Code and models will be released upon publication.
@@ C Morton-Order-Preserving Coordinate Transformation
A set of candidate parent coordinates is obtained by halving the integer coordinates,
@@
In practice, this occupancy-label generation uses only inexpensive integer operations on the child coordinates and avoids the $2{\times}2{\times}2$ down-convolution and kernel-map overhead common in conventional implementations.
@@ D Runtime Impact of Morton-Order Hierarchy
Fig. S3 provides a high-level comparison, at a single bit-depth level during compression, between explicit per-level sorting and our Morton-order-preserving hierarchy. In sparse-tensor-based geometry compression, it is essential that the encoder and decoder operate on exactly the same coordinate order so that occupancy bits are generated and consumed with aligned voxel indices. Existing approaches satisfy this by performing identical sorting at every downscaling and upsampling step in both encoding and decoding. In contrast, ELiC performs a single initial 3D Morton sort of the input voxel coordinates $\mathbf{C}^{(B)}$ and then applies downscaling and upsampling that preserve this global order, eliminating the per-level re-sorting required by prior methods in both encoding and decoding.
@@ E Asynchronous Arithmetic Encoding for Latency Reduction
In the ELiC pipeline, the neural network inference and arithmetic encoding operate sequentially, which introduces idle time on both CPU and GPU. To improve runtime efficiency, we implement an asynchronous pipeline that overlaps the neural network forward pass and arithmetic encoding on separate compute threads.
@@ G Rate-Distortion Curves
At the 16- and 15-bit coordinate inputs, the ELiC variants exhibit notably superior compression performance. However, at 13- and 12-bit coordinate inputs, a slight degradation is observed. This trend likely stems from the design of the training objective in Eq.( 10), which minimizes the total bit count for the entire hierarchy. As a result, the model tends to allocate more capacity to support the highly sparse upper bit-depths, where a larger number of bits are required. Future work should explore strategies to balance optimization across all bit-depth levels for more uniformly strong performance.

### v2

@@ Abstract
Hierarchical LiDAR geometry compression encodes voxel occupancies from low to high bit-depths, yet prior methods treat each depth independently and re-estimate local context from coordinates at every level, limiting compression efficiency. We present ELiC, a real-time framework that combines cross-bit-depth feature propagation, a Bag-of-Encoders (BoE) selection scheme, and a Morton-order-preserving hierarchy. Cross-bit-depth propagation reuses features extracted at denser, lower depths to support prediction at sparser, higher depths. BoE selects, per depth, the most suitable coding network from a small pool, adapting capacity to observed occupancy statistics without training a separate model for each level. The Morton hierarchy maintains global Z-order across depth transitions, eliminating per-level sorting and reducing latency. Together these components improve entropy modeling and computation efficiency, yielding state-of-the-art compression at real-time throughput on Ford and SemanticKITTI. Code and pretrained models are available at https://github.com/moolgom/ELiCv1.
@@ C Morton-Order-Preserving Coordinate Transformation
A duplicated set of parent coordinates is obtained by halving the integer coordinates,
@@
In practice, this occupancy-label generation uses only inexpensive integer operations on the child coordinates and avoids the conventional $2{\times}2{\times}2$ down-convolution with fixed kernel weights $[1,2,4,8,16,32,64,128]$ and the associated kernel-map construction overhead.
@@
The reference implementation of this Morton-order-preserving hierarchy and occupancy labeling procedure is available in the morton.py file in the code repository.
@@ D Runtime Impact of Morton-Order Hierarchy
Fig. S3 provides a high-level comparison, at a single bit-depth level during compression, between explicit per-level sorting and our Morton-order-preserving hierarchy. In sparse-tensor-based geometry compression, it is essential that the encoder and decoder operate on exactly the same coordinate order so that occupancy bits are generated and consumed with aligned voxel indices. Existing sparse-tensor-based approaches satisfy this by performing identical sorting at every downscaling and upsampling step in both encoding and decoding. In contrast, ELiC performs a single initial 3D Morton sort of the input voxel coordinates $\mathbf{C}^{(B)}$ and then applies downscaling and upsampling that preserve this global order, eliminating the per-level re-sorting required by prior methods in both encoding and decoding.
@@
All runtimes reported in this supplementary material were measured on a machine equipped with an NVIDIA GeForce RTX 3090 GPU and an Intel Core i9-9900K CPU with 64 GB of RAM.
@@ E Asynchronous Arithmetic Encoding for Latency Reduction
In the ELiC pipeline, the neural network inference and arithmetic encoding operate sequentially. This means the CPU must wait idly for the GPU to finish the neural network forward pass before the arithmetic encoding can begin, and conversely, the GPU remains idle while the CPU performs the encoding operation. To improve runtime efficiency, we implement an asynchronous pipeline that overlaps these two tasks on separate compute threads.
@@
Note that the runtime numbers reported in the main paper do not include this additional speed-up from asynchronous arithmetic encoding.
@@ G Rate-Distortion Curves
At the 16- and 15-bit coordinate inputs, the ELiC variants exhibit notably superior compression performance. However, at 13- and 12-bit coordinate inputs, a slight degradation is observed. This trend likely stems from the design of the training objective in Eq. 10, which minimizes the total bit count for the entire hierarchy. As a result, the model tends to allocate more capacity to support the highly sparse upper bit-depths, where a larger number of bits is required. Future work should explore strategies to balance optimization across all bit-depth levels for more uniformly strong performance.